Technical Staff Blog

Category archives: Service Issues

Tstaff announcements related to general IT service issues

RSS feed of Service Issues

Last update on .

During the final weeks of the semester, the technical staff encountered issues with routine maintenance of our GPFS filesystem.  IBM suggested that there was corruption in the filesystem, and recommended that we take it offline to repair it.  We postponed that work until after commencement, to avoid disrupting the final weeks of classes, a conference ...

Last update on .

CIS has reported that this issue has been resolved. If you are still experiencing issues please contact them at the number below.

 

CIS has reported that a significant amount of users are experiencing an outage of Brown email (Gmail). More info from CIS below:

We are aware of a problem with Google Email - Brown Gmail ...

Last update on .

All critical CS Department services have been restored. An interruption in external services resulted in a cascading DNS failure and overall service instability, starting mid-morning on Thursday of this week.  After repeated efforts to diagnose and workaround the issue, we reached out to CIS who stuck with us throughout today to diagnose. Ultimately, we were ...

Last update on .

Late Friday evening into Saturday CIS expereinced a failure of one of the VMWare servers in the stack they provide to us. A number of our internal services depend on the virtual machines hosted by that server, which resulted in intermittent issues through the department (website, dhcp, vpn, ssh, and others).

CIS has temporarily migrated ...

Last update on .

TStaff has migrated all print functions from the old print host to a new and improved print host to alleviate and resolve the many print issues that existed on the old print server. Due to some driver changes and a newer version of CUPS some user intervention may be needed. Users experiencing issues with printing ...

Last update on .

On 12/1, from approximately 10PM to 12AM, a department GPFS node became unresponsive, causing our FastX and SSH gateways to follow in turn. All services have returned to normal.

We apologize for any inconvenience. If you experience any continuing issues connecting to the department via either FastX or SSH, please contact us at problem ...

Last update on .

Due to campus-wide service outages, the main sections of the CS Dept website, including the home page, have been inaccessible for an extended period of time today, and remain down.  We are monitoring CIS' progress in addressing their issues, and we will ensure that the CS website returns to service when they are resolved.

Last update on .

Common SSH Issues/Fixes

ssh.png

On August 24th, the ssh.cs.brown.edu gateway servers were upgraded to stretch and bound to the CIS OIM domain, ad.brown.edu.  The sunlab and mlab ssh client systems were also upgrade on that day and bound to ad.brown.edu.  As of version 7.0, the openssh server has disabled support for ...